OpenStudy (anonymous):

Determine the a) x-intercept and b) y-intercept for the graph of the given equation 3x – 2y = 6 Take your given equation and solve for “y”. 3x – 2y = 6 please show steps thank you

OpenStudy (anonymous):

First solve for y. 3x - 2y = 6 -2y = -3x + 6 y = 2/3x - 3 The form is y = mx + b b = y-intercept. Therefore, -3 is the y-intercept. Then plug in 0 for y and solve because that's where the line crosses the x-axis. 0 = 2/3x - 3 3 = 2/3x 9/2 = x

OpenStudy (btaylor):

X intercept is when y=0. So plug 0 in for y, and solve for x. You get x=2. So x intercept is (2,0) Y intercept is when x=0. So plug 0 in for x, and solve for y. You get y= -3. So y intercept is (0,-3) slope-intercept form of a line: y=mx+b. You know b (the y intercept). Now, plug the other known point in for x and y, and you get 0 = 2m-3 solve for m, you get 2/3

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You're welcome.